OpenStudy (anonymous):

f(x) = x² + 2x - 3 find a zero of this function..show how u got that answer

OpenStudy (anonymous):

are you familiar with factoring?

OpenStudy (anonymous):

Yea like for examplt (1+3)(1-3)

OpenStudy (anonymous):

factoring f(x) would give you (x+3)(x-1).... i don't know where you got (1+3)(1-3) from

OpenStudy (anonymous):

i was just using it as an explample of factoring

OpenStudy (anonymous):

ok, well i factored it for you. that should give you the roots of f(x)

OpenStudy (anonymous):

Thank you so that means the the zero is -3 correct?

jhonyy9 (jhonyy9):

x^2 +2x -3 = 0 (x-1)(x+3) = 0 this mean x-1=0 x_1= 1 and x+3=0 x_2 =-3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yes thank you